OpenStudy (selena2345):

The mass of Earth is 5.97x10^24 kilograms. The mass of the moon is 7.34x10^22 kilograms. Part A: What is the combined mass of Earth and the Moon ? Express your in scientific notation. Part B: How many times as great is the mass of the Moon ? Express your answer in scientific notation.

OpenStudy (chaseffs):

5.97 x 10^24 + 7.34 x 10^22 = 5.97 x 10^24 + .0734 x 10^24 = 6.0434 x 10^24 5.97 x 10^24 / 7.34 x 10^22 = (5.97/7.34) x 10^2 = .8133 x 10^2 = 8.133 x 10^1

OpenStudy (selena2345):

is this Part A ? @ChaseFFS

OpenStudy (chaseffs):

Part A is up to. Part B is on the bottom.

OpenStudy (chaseffs):

Up top*

OpenStudy (selena2345):

Thanks Chase @ChaseFFS

OpenStudy (chaseffs):

No problem. :)